# A General Latent Embedding Approach for Modeling Non-uniform High-dimensional Sparse Hypergraphs with Multiplicity

### v1

@@ 3 Identifiability, estimation and statistical theory
or equivalently, $y_{ji}$’s are conditionally independent Bernoulli random variables with means $p_{ji}=\mathbb{P}(y_{ji}=1\mid\boldsymbol{\Xi}_{m,n}).$ We use the superscript ``"``"``*"` ` ∗ " to denote the true model parameters: $\boldsymbol{\Xi}_{m,n}^{*}=\big{\{}\beta_{m,n}^{*},\{\alpha_{i}^{*}\}_{i=1}^{n},\{\mathbf{z}_{i}^{*}\}_{i=1}^{n},\{\mathbf{f}_{j}^{*}\}_{j=1}^{m}\big{\}}$. Denote $\boldsymbol{\alpha}^{*}=(\alpha^{*}_{1},\cdots,\alpha_{n}^{*})^{\top}\in\mathbb{R}^{n}$, $\mathbf{F}^{*}=(\mathbf{f}_{1}^{*},\cdots,\mathbf{f}_{m}^{*})^{\top}\in\mathbb{R}^{m\times K}$, $\mathbf{Z}^{*}=(\mathbf{z}_{1}^{*},\cdots,\mathbf{z}_{n}^{*})^{\top}\in\mathbb{R}^{n\times K}$, and $\boldsymbol{\Theta}^{*}=(\theta_{ji}^{*})_{m\times n}=\beta^{*}_{m,n}\mathbf{1}_{m}\mathbf{1}_{n}^{\top}+\mathbf{1}_{m}\boldsymbol{\alpha}^{*\top}+\mathbf{F}^{*}\mathbf{Z}^{*\top}\in\mathbb{R}^{m\times n}$. In the following, Section 3.1 considers the identifiability of $\beta^{*}_{m,n}$ and $\boldsymbol{\alpha}^{*}$. A constrained maximum likelihood estimator is studied and its $F$-consistency in the estimation for $\boldsymbol{\Theta}^{*}$, $\boldsymbol{\alpha}^{*}$ and $\beta^{*}_{m,n}$ is derived. Section 3.2 further considers the identification and estimation of all the individual components in $\boldsymbol{\Xi}_{m,n}^{*}$. Another constrained maximum likelihood estimator with additional constraints is studied. We show entry-wise consistency of the estimators and derive their asymptotic distributions. These results further enable consistent estimation and asymptotic inference for each $\theta^{*}_{ji}$ and $p_{ji}^{*}=\sigma(\theta^{*}_{ji})$.
@@
[…]
@@
where $C_{1},C_{2}>0$ are some constants, and $\mathbf{J}_{m}=\mathbf{I}_{m}-m^{-1}\mathbf{1}_{m}\mathbf{1}_{m}^{\top}$ is the centering operator. The boundedness constraint $\max_{i,j}\{|\alpha_{i}|,\|\mathbf{z}_{i}\|_{2},\|\mathbf{f}_{j}\|_{2}\}\leq C_{1}$ is similar to those employed in parameter estimation for fitting dyadic network latent space models (e.g., Ma et al., 2020). Note that we allow $\beta_{m,n}^{*}$ to diverge towards $\infty$- ∞ as $m,n\to\infty$, thus the parameter $C_{\beta_{m,n}}$ serves to control the rate of divergence of $\beta$ in the estimation, and the choice of the rate will be discussed in Remark 3.3 below.
@@
Let $\hat{\boldsymbol{\Theta}}=\hat{\beta}\mathbf{1}_{m}\mathbf{1}_{n}^{\top}+\mathbf{1}_{m}\hat{\boldsymbol{\alpha}}^{\top}+\hat{\mathbf{F}}\hat{\mathbf{Z}}^{\top}$ where $(\hat{\beta},\hat{\boldsymbol{\alpha}},\hat{\mathbf{F}},\hat{\mathbf{Z}})$ is the solution to (3.2). For any $\epsilon>0$, as long as $\exp(\beta^{*}_{m,n})\gtrsim\log\big{(}(m\vee n)/\epsilon\big{)}/(m\wedge n)$ and $(\beta_{m,n}^{*},\boldsymbol{\alpha}^{*},\mathbf{F}^{*},\mathbf{Z}^{*})\in{\cal F}_{1}(C_{\beta_{m,n}})$, we have
@@
Suppose that there exists $s_{m,n}\to\infty$ such that $\exp(\beta_{m,n}^{*})\gg s_{m,n}/(mn)$ as $m,n\to\infty$. Then the choice of $C_{\beta_{m,n}}$ in Remark 3.3 satisfies $C_{\beta_{m,n}}\asymp-\beta^{*}_{m,n}$ with probability approaching $1$ as $m,n\to\infty$. Consequently, there exist constants $C^{\prime}\geq 1$ in Remark 3.3 and $C_{1},C_{2}>0$ in (3.1) such that
$\mathbb{P}\big{\{}(\beta_{m,n}^{*},\boldsymbol{\alpha}^{*},\mathbf{F}^{*},\mathbf{Z}^{*})\in{\cal F}_{1}(C_{\beta_{m,n}})\big{\}}\to 1$
as $m,n\to\infty$.
@@
Let $(\hat{\beta},\hat{\boldsymbol{\alpha}},\hat{\mathbf{F}},\hat{\mathbf{Z}})$ be the solution to (3.2). For any $\epsilon>0$, as long as $\exp(\beta^{*}_{m,n})\gtrsim\log\big{(}(m\vee n)/\epsilon\big{)}/(m\wedge n)$ and $(\beta_{m,n}^{*},\boldsymbol{\alpha}^{*},\mathbf{F}^{*},\mathbf{Z}^{*})\in{\cal F}_{1}(C_{\beta_{m,n}})$, we have
@@ 3.2 Entry-wise consistent estimation and asymptotic inference for embedding vectors
which preserves the likelihood in (1.1), making individual embeddings not identifiable from observed data. To address the identifiability issue, we assume a set of embedding vector constraints commonly employed as follows:
[…]
@@
[…]
@@
For the technical derivation of entry-wise consistent results and asymptotic distributions, we further set $C_{3}=-C_{3}^{\prime}\cdot C_{\beta_{m,n}}$ for some $C_{3}^{\prime}\in(0,1)$. As shown in the following proposition, this choice of $C_{3}$ retains the good properties of ${\cal F}_{1}(C_{\beta_{m,n}})$, where $C_{\beta_{m,n}}$ is taken as in Remark 3.3. The additional constraint for $C_{3}$ is mainly for the technical reason, and practically in algorithmic implementation, we find that the choice of $C_{3}$ can be flexible and does not significantly change the results.
@@
Following the conditions in Proposition 3.1, there exist constants $C^{\prime}\geq 1$ in Remark 3.3 and $C_{3}^{\prime}\in(0,1),C_{4},C_{5}>0$ such that
$\mathbb{P}\big{\{}(\boldsymbol{\alpha}^{\dagger},\mathbf{f}^{*},\mathbf{z}^{*})\in{\cal F}_{2}(C_{\beta_{m,n}})\big{\}}\to 1$
as $m,n\to\infty$.
@@
[…]
@@
Let $(\hat{\boldsymbol{\alpha}},\hat{\mathbf{z}},\hat{\mathbf{f}})$ be the solution to (3.6). Under Condition 3.1, assume that $(\boldsymbol{\alpha}^{\dagger},\mathbf{z}^{*},\mathbf{f}^{*})\in{\cal F}_{2}(C_{\beta_{m,n}})$. Define
@@
with $\mbox{with }{\cal P}(\mathbf{z},\mathbf{f}|\lambda_{m,n})=$ $~{}\frac{\lambda_{m,n}mn}{8}\Big{\|}\text{diag}\Big{(}\frac{1}{n}\mathbf{Z}^{\top}\mathbf{Z}-\frac{1}{m}\mathbf{F}^{\top}\mathbf{F}\Big{)}\Big{\|}_{F}^{2}+\frac{\lambda_{m,n}mn}{2}\Big{\|}\frac{\mathbf{F}^{\top}\mathbf{1}_{m}}{m}\Big{\|}_{2}^{2}$
@@
Since the solution remains unchanged for any $\lambda_{m,n}>0$, we can take advantage of the penalization term ${\cal P}(\mathbf{z},\mathbf{f}|\lambda_{m,n})$ in (3.9) when analyzing (3.6): in combination with the Hessian matrix of the log-likelihood function, the Hessian of the penalty function ${\cal P}$ expands the entire linear space $\mathbb{R}^{n(K+1)+mK}$ with high probability. Similar Lagrangian multiplier strategies have been used in the literature, such as Silvey (1959) and El-Helbawy & Hassan (1994) for Lagrangian multiplier test and Wang (2022) for generalized factor models with application to factor-augmented regressions. Due to space constraints, detailed theoretical analyses are in the supplementary material and we summarize the main results below. In Theorem 3.2, we establish entry-wise estimation error bounds for (3.6). We first introduce two conditions for achieving the entry-wise consistency results: one on the requirement for the order-adjusting parameter, and the other on the constraint parameters in (3.6).
@@
Theorem 3.2 (Entry-wise consistent estimation of the embedding vectors and un-centered vertex degrees).
Let $(\hat{\boldsymbol{\alpha}},\hat{\mathbf{z}},\hat{\mathbf{f}})$ be the solution to (3.6). Under Conditions 3.1, 3.2 and 3.3, assume that $(\boldsymbol{\alpha}^{\dagger},\mathbf{z}^{*},\mathbf{f}^{*})\in{\cal F}_{2}(C_{\beta_{m,n}})$. Then we have
@@
for any $\epsilon>0$, as $(m,n)\to\infty$.
Let $\mathbf{P}^{*}=(p^{*}_{ji})_{m\times n}=\sigma(\boldsymbol{\Theta}^{*})$. The next corollary demonstrates entry-wise consistent estimation of $\theta^{*}_{ji}$ and $p_{ji}^{*}$ based on (3.6).
Corollary 3.2 (Entry-wise consistent $\boldsymbol{\Theta}$ and $\mathbf{P}$ estimation).
Let $(\hat{\boldsymbol{\alpha}},\hat{\mathbf{z}},\hat{\mathbf{f}})$ be the solution to (3.6), $\hat{\boldsymbol{\Theta}}=(\hat{\theta}_{ji})_{m\times n}$ with $\hat{\theta}_{ji}=\hat{\alpha}_{i}+\hat{\mathbf{f}}_{j}^{\top}\hat{\mathbf{z}}_{i}$, and $\hat{\mathbf{P}}=\sigma(\hat{\boldsymbol{\Theta}})$. Then under the conditions in Theorem 3.2, for any $\epsilon>0$, we have
@@
In the following, we present asymptotic distributions for the estimators in (3.6). Define $\sigma^{\prime}(x)=\exp(x)/\{1+\exp(x)\}^{2}$ as the derivative of $\sigma(x)$. For ease of notation, we let $\mathbf{q}_{j}^{*}=(1,\mathbf{f}_{j}^{*\top})^{\top}$ and $\boldsymbol{\nu}_{i}=(\alpha_{i},\mathbf{z}_{i}^{\top})^{\top}$. Specify a pair $(j,i)\in[m]\times[n]$. We define
$\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},m,n)}=\bigg{(}\frac{1}{m}\sum_{j=1}^{m}\Big{[}\sigma^{\prime}(\theta^{*}_{ji})\exp(-\beta_{m,n}^{*})\mathbf{q}_{j}^{*}\mathbf{q}_{j}^{*\top}\Big{]}\bigg{)}^{-1}~{}\text{ and }$ (3.10)
$\boldsymbol{\Sigma}_{(\mathbf{f}_{j},m,n)}=\bigg{(}\frac{1}{n}\sum_{i=1}^{n}\Big{[}\sigma^{\prime}(\theta^{*}_{ji})\exp(-\beta_{m,n}^{*})\mathbf{z}_{i}^{*}\mathbf{z}_{i}^{*\top}\Big{]}\bigg{)}^{-1}.$
These definitions are strategically chosen to ensure that the magnitudes of $\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},m,n)}$ and $\boldsymbol{\Sigma}_{(\mathbf{f}_{j},m,n)}$ maintain a constant order as $m,n\to\infty$. This strategy is consistently applied in the remaining presentation of the asymptotic distribution results. The intention is to keep the “$\boldsymbol{\Sigma}$” terms and the ”$\mathrm{Var}$” terms that will be introduced later at a constant order, making it easier to see the rate of the asymptotic variance-covariance. Let $\boldsymbol{\nu}^{*}_{i}=(\alpha_{i}^{\dagger},\mathbf{z}_{i}^{*\top})^{\top}$ for $i\in[n]$. The following theorem presents asymptotic distributions for $(\hat{\alpha}_{i},\hat{\mathbf{z}}_{i}^{\top})^{\top}$ and $\hat{\mathbf{f}}_{j}$.
Theorem 3.3 (Individual asymptotic distributions).
Let $(\hat{\boldsymbol{\alpha}},\hat{\mathbf{z}},\hat{\mathbf{f}})$ be the solution to (3.6) and $\hat{\boldsymbol{\nu}}_{i}=(\hat{\alpha}_{i},\hat{\mathbf{z}}_{i}^{\top})^{\top}$. Under the conditions in Theorem 3.2, as $m,n\to\infty$, we have
@@
Furthermore, let
${\boldsymbol{\Omega}}_{ji}^{(m,n)}=\begin{pmatrix}\frac{1}{m}\sum_{j=1}^{m}\Big{[}\sigma^{\prime}(\theta^{*}_{ji})\exp(-\beta_{m,n}^{*})\mathbf{q}_{j}^{*}\mathbf{q}_{j}^{*\top}\Big{]}&\frac{1}{m^{1/2}n^{1/2}}\sigma^{\prime}(\theta^{*}_{ji})\exp(-\beta_{m,n}^{*})\mathbf{q}_{j}^{*}\mathbf{z}_{i}^{*\top}\\ \frac{1}{m^{1/2}n^{1/2}}\sigma^{\prime}(\theta^{*}_{ji})\exp(-\beta_{m,n}^{*})\mathbf{z}_{i}^{*}\mathbf{q}_{j}^{*\top}&\frac{1}{n}\sum_{i=1}^{n}\Big{[}\sigma^{\prime}(\theta^{*}_{ji})\exp(-\beta_{m,n}^{*})\mathbf{z}_{i}^{*}\mathbf{z}_{i}^{*\top}\Big{]}\end{pmatrix},$
and define
$\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},\mathbf{f}_{j},m,n)}=\begin{bmatrix}\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},m,n)}&\mathbf{0}\\ \mathbf{0}&\boldsymbol{\Sigma}_{(\mathbf{f}_{j},m,n)}\end{bmatrix}{\boldsymbol{\Omega}}_{ji}^{(m,n)}\begin{bmatrix}\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},m,n)}&\mathbf{0}\\ \mathbf{0}&\boldsymbol{\Sigma}_{(\mathbf{f}_{j},m,n)}\end{bmatrix}.$
Then Theorem 3.4 obtains the joint asymptotic distribution for $(\hat{\alpha}_{i},\hat{\mathbf{z}}_{i},\hat{\mathbf{f}}_{j})$.
Theorem 3.4 (Joint asymptotic distributions).
Let $(\hat{\boldsymbol{\alpha}},\hat{\mathbf{z}},\hat{\mathbf{f}})$ be the solution to (3.6) and $\hat{\boldsymbol{\nu}}_{i}=(\hat{\alpha}_{i},\hat{\mathbf{z}}_{i}^{\top})^{\top}$. Under the conditions in Theorem 3.2, and further assume that $\exp(\beta_{m,n}^{*}/2)\gg(m^{\frac{1}{2}}+n^{\frac{1}{2}})(m+n)^{\epsilon^{*}}/(m\wedge n)$, then we have
@@
as $m,n\to\infty$.
Note that as $m,n\to\infty$, $\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},\mathbf{f}_{j},m,n)}$ becomes block-wise diagonal, implying that $\hat{\boldsymbol{\nu}}_{i}$ and $\hat{\mathbf{f}_{j}}$ are asymptotically independent. Building upon the joint limit distributions in Theorem 3.4, we present limit distributions for estimates of $\theta_{ji}^{*}$ and $p_{ji}^{*}$ in Corollary 3.3.
Let
$\mathrm{Var}_{m,n}{(\theta_{ji})}=(m\wedge n)\begin{pmatrix}\mathbf{q}_{j}^{*}\\ \mathbf{z}_{i}^{*}\end{pmatrix}^{\top}\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},\mathbf{f}_{j},m,n)}\begin{pmatrix}\mathbf{q}_{j}^{*}\\ \mathbf{z}_{i}^{*}\end{pmatrix}~{}\text{ and}$
$\mathrm{Var}_{m,n}$ ${(p_{ji})}=(m\wedge n)(\sigma^{\prime\prime}(\theta_{ji}^{*})\exp(-\beta_{m,n}^{*}))^{2}\begin{pmatrix}\mathbf{q}_{j}^{*}\\ \mathbf{z}_{i}^{*}\end{pmatrix}^{\top}\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},\mathbf{f}_{j},m,n)}\begin{pmatrix}\mathbf{q}_{j}^{*}\\ \mathbf{z}_{i}^{*}\end{pmatrix}.$
Let $(\hat{\boldsymbol{\alpha}},\hat{\mathbf{z}},\hat{\mathbf{f}})$ be the solution to (3.6). Under the conditions in Theorem 3.4, let $\hat{\theta}_{ji}=\hat{\alpha}_{i}+\hat{\mathbf{f}}_{j}^{\top}\hat{\mathbf{z}}_{i}$ and $\hat{p}_{ji}=\sigma(\hat{\theta}_{ji})$, then we have
[…]
Theorem 3.3 and Corollary 3.3 provide the basis for constructing confidence intervals for $(\alpha_{i}^{\dagger},\mathbf{z}_{i}^{*},\mathbf{f}_{j}^{*})$, $\theta_{ji}^{*}$ and $p_{ji}^{*}$. To complete the inference task, we plug in the estimators $(\hat{\boldsymbol{\alpha}},\hat{\mathbf{z}},\hat{\mathbf{f}})$ into the covariance expressions. Note that $\beta_{m,n}^{*}$ does not need to be estimated and plugged in for inference implementation, as it cancels out in the covariance terms. We plug in $(\hat{\theta}_{ji},\hat{\mathbf{z}}_{i},\hat{\mathbf{f}}_{j})$ for $(\theta_{ji}^{*},\mathbf{z}_{i}^{*},\mathbf{f}_{j}^{*})$ in (3.10) and ${\boldsymbol{\Omega}}_{ji}^{(m,n)}$. Let $\widehat{\boldsymbol{\Sigma}}_{(\boldsymbol{\nu}_{i},m,n)}$ and $\widehat{\boldsymbol{\Sigma}}_{(\mathbf{f}_{j},m,n)}$ denote the plug-in estimators for ${\boldsymbol{\Sigma}}_{(\boldsymbol{\nu}_{i},m,n)}$ and ${\boldsymbol{\Sigma}}_{(\mathbf{f}_{j},m,n)}$, respectively. Let $\widehat{\mathrm{Var}}_{m,n}{(\theta_{ji})}$ and $\widehat{\mathrm{Var}}_{m,n}{(p_{ji})}$ be the plug-in estimators for ${\mathrm{Var}}_{m,n}{(\theta_{ji})}$ and ${\mathrm{Var}}_{m,n}{(p_{ji})}$, respectively. The following corollary establishes the consistency of the plug-in estimators.
@@
Combining Theorem 3.3, Corollaries 3.3 and 3.4 and applying Slutsky’s Theorem, we have that the asymptotic normality results hold for the plug-in variance-covariance.

### v2

@@ 3 Identifiability, estimation and statistical theory
or equivalently, $y_{ji}$’s are conditionally independent Bernoulli random variables with means $p_{ji}=\mathbb{P}(y_{ji}=1\mid\boldsymbol{\Xi}_{m,n}).$ We use the superscript ``*" to denote the true model parameters: $\boldsymbol{\Xi}_{m,n}^{*}=\big\{\beta_{m,n}^{*},\{\alpha_{i}^{*}\}_{i=1}^{n},\{\mathbf{z}_{i}^{*}\}_{i=1}^{n},\{\mathbf{f}_{j}^{*}\}_{j=1}^{m}\big\}$. Denote $\boldsymbol{\alpha}^{*}=(\alpha^{*}_{1},\cdots,\alpha_{n}^{*})^{\top}\in\mathbb{R}^{n}$, $\mathbf{F}^{*}=(\mathbf{f}_{1}^{*},\cdots,\mathbf{f}_{m}^{*})^{\top}\in\mathbb{R}^{m\times K}$, $\mathbf{Z}^{*}=(\mathbf{z}_{1}^{*},\cdots,\mathbf{z}_{n}^{*})^{\top}\in\mathbb{R}^{n\times K}$, and $\boldsymbol{\Theta}^{*}=(\theta_{ji}^{*})_{m\times n}=\beta^{*}_{m,n}\mathbf{1}_{m}\mathbf{1}_{n}^{\top}+\mathbf{1}_{m}\boldsymbol{\alpha}^{*\top}+\mathbf{F}^{*}\mathbf{Z}^{*\top}\in\mathbb{R}^{m\times n}$.
In the following, Section 3.1 examines the identifiability of $\beta^{*}_{m,n}$ and $\boldsymbol{\alpha}^{*}$. We study a constrained maximum likelihood estimator and establish its $F$-consistency in estimating $\boldsymbol{\Theta}^{*}$, $\boldsymbol{\alpha}^{*}$, and $\beta^{*}_{m,n}$. Section 3.2 further addresses the identification and estimation of all individual components in $\boldsymbol{\Xi}_{m,n}^{*}$. We study another constrained maximum likelihood estimator with a more comprehensive set of constraints, show entry-wise consistency of the estimators, and derive their asymptotic distributions. These results further enable consistent estimation and asymptotic inference for each $\theta^{*}_{ji}$ and $p_{ji}^{*}=\sigma(\theta^{*}_{ji})$, for all $(j,i)\in[m]\times[n]$.
@@
${\cal F}_{1}(C_{\beta_{m,n}})$$\big\{$ $(\beta,\boldsymbol{\alpha},\mathbf{F},\mathbf{Z})\in\mathbb{R}\times\mathbb{R}^{n}\times\mathbb{R}^{m\times K}\times\mathbb{R}^{n\times K}:\mathbf{J}_{m}\mathbf{F}=\mathbf{F},\mathbf{1}_{n}^{\top}\boldsymbol{\alpha}=0,$ (3.1)
@@
where $C_{1},C_{2}>0$ are some constants, and $\mathbf{J}_{m}=\mathbf{I}_{m}-m^{-1}\mathbf{1}_{m}\mathbf{1}_{m}^{\top}$ is the centering operator. The boundedness constraint $\max_{i,j}\{|\alpha_{i}|,\|\mathbf{z}_{i}\|_{2},\|\mathbf{f}_{j}\|_{2}\}\leq C_{1}$ is similar to those employed in parameter estimation for fitting dyadic network latent space models (e.g., Ma et al., 2020). Note that we allow $\beta_{m,n}^{*}$ to diverge towards -$\infty$ as $m,n\to\infty$, thus the parameter $C_{\beta_{m,n}}$ serves to control the rate of divergence of $\beta$ in the estimation, and the choice of the rate will be discussed in Remark 3.3 below.
@@
[…]
@@
Suppose that there exists $s_{m,n}\to\infty$ such that $\exp(\beta_{m,n}^{*})\gg s_{m,n}/(mn)$ as $m,n\to\infty$. Then the choice of $C_{\beta_{m,n}}$ in Remark 3.3 satisfies $C_{\beta_{m,n}}\asymp-\beta^{*}_{m,n}$ with probability approaching 11 as $m,n\to\infty$. Consequently, there exist constants $C^{\prime}\geq 1$ in Remark 3.3 and $C_{1},C_{2}>0$ in (3.1) such that $\mathbb{P}\big\{(\beta_{m,n}^{*},\boldsymbol{\alpha}^{*},\mathbf{F}^{*},\mathbf{Z}^{*})\in{\cal F}_{1}(C_{\beta_{m,n}})\big\}\to 1$ as $m,n\to\infty$.
@@
[…]
@@ 3.2 Entry-wise consistent estimation and asymptotic inference for embedding vectors
which preserves the likelihood in (1.1), making individual embeddings not identifiable from observed data. To address the identifiability issue, we assume a set of embedding vector constraints as follows:
$\text{(C1) Both}~\mathbf{F}^{*\top}\mathbf{F}^{*}~\text{and}~\mathbf{Z}^{*\top}\mathbf{Z}^{*}~\text{are diagonal;}~~\text{(C2)}~\frac{\mathbf{F}^{*\top}\mathbf{F}^{*}}{m}=\frac{\mathbf{Z}^{*\top}\mathbf{Z}^{*}}{n};~~\text{(C3)}~\mathbf{J}_{m}\mathbf{F}^{*}=\mathbf{F}^{*}.$
@@
${\cal F}_{2}(C_{\beta_{m,n}})$$\{$ $(\boldsymbol{\alpha},\mathbf{z},\mathbf{f})\in\mathbb{R}^{n}\times\mathbb{R}^{nk}\times\mathbb{R}^{mk}:-C_{\beta_{m,n}}\leq\sum_{i=1}^{n}\alpha_{i}/n\leq C_{3},$ (3.5)
@@
For the technical derivation of entry-wise consistent results and asymptotic distributions, we further set $C_{3}=-C_{3}^{\prime}\cdot C_{\beta_{m,n}}$ for some $C_{3}^{\prime}\in(0,1)$. As shown in the following proposition, this choice of $C_{3}$ retains the good properties of ${\cal F}_{1}(C_{\beta_{m,n}})$, where $C_{\beta_{m,n}}$ is taken as in Remark 3.3. The additional constraint for $C_{3}$ is mainly for technical purposes, and practically in algorithmic implementation, we find that the choice of $C_{3}$ can be flexible and does not significantly change the results.
@@
Following the conditions in Proposition 3.1, there exist constants $C^{\prime}\geq 1$ in Remark 3.3 and $C_{3}^{\prime}\in(0,1),C_{4},C_{5}>0$ such that $\mathbb{P}\big\{(\boldsymbol{\alpha}^{\dagger},\mathbf{f}^{*},\mathbf{z}^{*})\in{\cal F}_{2}(C_{\beta_{m,n}})\big\}\to 1$ as $m,n\to\infty$.
@@
$\max_{\boldsymbol{\alpha},\mathbf{z},\mathbf{f}}{\cal L}({\mathbf{Y}}|\boldsymbol{\alpha},\mathbf{z},\mathbf{f})~~\text{s.t.}$ $\frac{1}{n}\mathbf{Z}^{\top}\mathbf{Z}=\frac{1}{m}\mathbf{F}^{\top}\mathbf{F},~\mathbf{Z}^{\top}\mathbf{Z}\text{ and }\mathbf{F}^{\top}\mathbf{F}\text{ are diagonal},$ (3.6)
@@
[…]
@@
$\mbox{with }{\cal P}(\mathbf{z},\mathbf{f}|\lambda_{m,n})=$ $~\frac{\lambda_{m,n}mn}{8}\Big\|\text{diag}\Big(\frac{1}{n}\mathbf{Z}^{\top}\mathbf{Z}-\frac{1}{m}\mathbf{F}^{\top}\mathbf{F}\Big)\Big\|_{F}^{2}+\frac{\lambda_{m,n}mn}{2}\Big\|\frac{\mathbf{F}^{\top}\mathbf{1}_{m}}{m}\Big\|_{2}^{2}$
@@
Since the solution remains unchanged for any $\lambda_{m,n}>0$, we can take advantage of the penalization term ${\cal P}(\mathbf{z},\mathbf{f}|\lambda_{m,n})$ in (3.9) when analyzing (3.6): in combination with the Hessian matrix of the log-likelihood function, the Hessian of the penalty function $\cal P$} expands the entire linear space $\mathbb{R}^{n(K+1)+mK}$ with high probability. Similar Lagrangian multiplier strategies have been used in the literature, such as Silvey (1959) and El-Helbawy & Hassan (1994) for Lagrangian multiplier test and Wang (2022) for generalized factor models with application to factor-augmented regressions. Due to space constraints, detailed theoretical analyses are in the supplementary material and we summarize the main results below.
We begin by introducing the following two conditions.
@@
Theorem 3.2 (Consistency of the embedding vectors and un-centered vertex degrees estimation).
[…]
@@
for any $\epsilon>0$, as $(m,n)\to\infty$. Moreover, we have an improved average consistency rate as
$\frac{1}{m}\sum_{j=1}^{m}\|\hat{\mathbf{f}}_{j}-\mathbf{f}_{j}^{*}\|_{2}^{2}+\frac{1}{n}\sum_{i=1}^{n}\|\hat{\mathbf{z}}_{i}-\mathbf{z}_{i}^{*}\|_{2}^{2}+\frac{1}{n}\|\hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}^{\dagger}\|_{2}^{2}=O_{p}\Big\{\frac{1}{(m\wedge n)\exp(\beta_{m,n}^{*})}\Big\}.$} (3.10)
[…]
Corollary 3.2 (Consistency of $\boldsymbol{\Theta}$ and $\mathbf{P}$ estimation).
[…]
@@
In the following, we present asymptotic distributions for the estimators in (3.6). Let $\boldsymbol{\nu}^{*}_{i}=(\alpha_{i}^{\dagger},\mathbf{z}_{i}^{*\top})^{\top}$ for $i\in[n]$. The following theorem presents the individual and joint asymptotic distributions of $(\hat{\alpha}_{i},\hat{\mathbf{z}}_{i}^{\top})^{\top}$ and $\hat{\mathbf{f}}_{j}$.
Theorem 3.3 (Latent embedding asymptotic distributions).
@@
  • (Individual) $\exp(\beta_{m,n}^{*}/2)m^{\frac{1}{2}}\boldsymbol{\Sigma}^{-\frac{1}{2}}_{(\boldsymbol{\nu}_{i},m,n)}(\hat{\boldsymbol{\nu}}_{i}-\boldsymbol{\nu}_{i}^{*})\overset{d}{\rightarrow}{\cal N}\big(\mathbf{0},\mathbf{I}\big)$ and
    $\exp(\beta_{m,n}^{*}/2)n^{\frac{1}{2}}\boldsymbol{\Sigma}_{(\mathbf{f}_{j},m,n)}^{-\frac{1}{2}}\big(\hat{\mathbf{f}}_{j}-\mathbf{f}_{j}^{*}\big)\overset{d}{\rightarrow}{\cal N}\big(\mathbf{0},\mathbf{I}\big)$;
  • (Joint) $\exp(\beta_{m,n}^{*}/2)\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},\mathbf{f}_{j},m,n)}^{-\frac{1}{2}}\big\{m^{\frac{1}{2}}(\hat{\boldsymbol{\nu}}_{i}-\boldsymbol{\nu}_{i}^{*})^{\top},n^{\frac{1}{2}}(\hat{\mathbf{f}}_{j}-\mathbf{f}_{j}^{*})^{\top}\big\}^{\top}\overset{d}{\longrightarrow}{\cal N}(\mathbf{0}_{2K+1},\mathbf{I}_{2K+1})$,
as $m,n\to\infty$, where $\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},m,n)}$, $\boldsymbol{\Sigma}_{(\mathbf{f}_{j},m,n)}$, and $\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},\mathbf{f}_{j},m,n)}$ denote the asymptotic variance terms, with their full definitions provided in Section 4.3.1 of the supplementary material.
Building upon the joint limit distributions in Theorem 3.3, we present limit distributions for estimates of $\theta_{ji}^{*}$ and $p_{ji}^{*}$ in Corollary 3.3.
@@
$\exp(\beta_{m,n}^{*}/2)(m\wedge n)^{\frac{1}{2}}\mathrm{Var}_{m,n}^{-\frac{1}{2}}{(\theta_{ji})}(\hat{\theta}_{ji}-\theta_{ji}^{*})\overset{d}{\longrightarrow}{\cal N}(0,1),$ (3.11)
$\exp(-\beta_{m,n}^{*}/$ $2)(m\wedge n)^{\frac{1}{2}}\mathrm{Var}^{-\frac{1}{2}}_{m,n}{(p_{ji})}(\hat{p}_{ji}-p_{ji}^{*})\overset{d}{\longrightarrow}{\cal N}(0,1),$ (3.12)
as $m,n\to\infty$, where $\mathrm{Var}_{m,n}{(\theta_{ji})}$ and $\mathrm{Var}_{m,n}{(p_{ji})}$ are asymptotic variance terms with their full definitions provided in Section 4.3.1 of the supplementary material.
@@
Combining Theorem 3.3, Corollaries 3.3 and 3.4 and applying Slutsky’s Theorem, we have that the asymptotic normality results hold for the plug-in variance-covariance terms.